North Company designs and manufactures machines that facilitate DNA sequencing. Depending on the intended purpose of each machine and its functions, each machine is likely to be unique. The job costing system in its Norfolk plant has five activity cost pools, in addition to direct materials and direct labor. Job TPY—2306 requires 1,000 printed-circuit boards. The cost per board that passes the final inspection is $240. On average, only 50% of the completed units pass the final inspection. The prime costs per completed board include $26 for direct materials and $5 for direct labor.

Required:

Information pertaining to manufacturing overhead for printed-circuit boards follows: Fill in the missing amount. (Round your answers to 2 decimal places.)

Activity Cost Driver Activity Driver Rate Activity Consumption per Board Factory Overhead per Board
Axial insertion Number of axial insertions $0.33 24.00
Hardware insertion Number of hardware insertions 4.00 54.80
Hand load Boothroyd time 5.00 38.55
Masking Number of points masked 0.16 90.00
Final test Test time 10.00 6.00
Answer #1

Answers

Activity Cost Driver Activity Driver Rate Activity Consumption per Board Factory Overhead per Board
Axial insertion Number of axial insertions $                            0.33 24 $                        7.92
Hardware insertion Number of hardware insertions $                            4.00 13.7 $                      54.80
Hand load Boothroyd time $                            7.71 5 $                      38.55
Masking Number of points masked $                            0.16 90 $                      14.40
Final test Test time $                            0.60 10 $                        6.00

--Workings

Activity Cost Driver Activity Driver Rate Activity Consumption per Board Factory Overhead per Board
Axial insertion Number of axial insertions 0.33 24 =24*0.33
Hardware insertion Number of hardware insertions 4 =54.8/4 54.8
Hand load Boothroyd time =38.55/5 5 38.55
Masking Number of points masked 0.16 90 =90*0.16
Final test Test time =6/10 10 6